# Pricing AI Model Accuracy

### v2

@@
Keywords: AI model market, model error trade-off, heterogeneous cost sensitivity, Hotelling model, firm specialization.
@@ 1 Introduction
While this goal is theoretically optimal, achieving a minimal error rate can be both costly and impractical for firms (Zhang et al., 2019). Improving multiple error dimensions simultaneously can be expensive and inefficient. In addition, an improvement in one dimension may lead to a worsening in another dimension, inducing an inverse trade-off across dimensions (Yaniv and Foster, 1995). A typical example of this is binary classification, where directly minimizing both Type I and Type II errors with a fixed sample size is typically infeasible (Mudge et al., 2012). Assuming a fixed total level of error, designing a model with fewer false positive (FP) errors would shrink the rejection region but also simultaneously lead to an increase in false negative (FN) error risk, and vice versa. As a result, model designers have to prioritize between FP and FN errors based on their expectations about consumer preferences. A standard statistical approach is fixing one dimension and optimizing the other (Maier and Lakens, 2022).
On the demand side, consumers often care less about the total model error and more about the specific types of errors a model makes, as some errors may be more costly to them than others. The bias-variance trade-off is an example of this inverse relation in model performance; lowering bias by improving model complexity typically leads to higher variance (Geman et al., 1992). One specific measure of model performance that consumers often consider is based on FP and FN rates (Cappelen et al., 2023). Different consumers exhibit varying levels of sensitivities to FP and FN rates, depending on their specific use cases and their associated cost sensitivities to FP and FN errors.
A prevalent shortcoming with generative AI models is hallucination, where the model generates entirely incorrect outputs. Hallucination rates, or more broadly FP rates, of the most popular models like ChatGPT remain quite significant (Rawte et al., 2023). Users employing AI for creative writing may have a completely different tolerance for hallucination compared to someone using the same model for disease diagnosis. This consumer heterogeneity can be captured by different sensitivities toward hallucination rates. For example, a scientific user may prefer a non-informative model output (i.e., “I don’t know”) as opposed to a model hallucination.
@@ 2 Related Literature
Our work builds upon past research at the intersection of machine learning and strategic behavior, focusing on how competition affects efficiency and the resulting equilibria. Previous literature has extensively analyzed competition in such settings. For example, Immorlica et al. (2011) propose an optimization-based framework for strategic competition, but does not include prices. Feng et al. (2022) examine the bias-variance trade-off in competitive environments, demonstrating a tendency for firms to favor variance-based errors over bias-based errors. Similarly, Jagadeesan et al. (2024) propose a classification model under which competition can potentially lead to worse overall model accuracy. Other studies adopt more information-driven approaches to competition (Liang, 2020) and explore coordination among firms to learn about consumer types (Gradwohl and Tennenholtz, 2023).
This paper differs from the related literature along multiple dimensions. First, we focus explicitly on competition over false positive (FP) and false negative (FN) rates rather than on the bias-variance trade-off. While the bias-variance framework is instrumental for understanding model complexity and generalization (Feng et al., 2022), the FP/FN trade-off is more specific to classification tasks. Jagadeesan et al. (2024) examines competition based on accuracy measured with Bayes risk, but firms do not choose prices. Classification problems are central to many real-world applications, such as medical diagnostics, fraud detection, and content filtering, where FP and FN errors carry distinct and significant costs. By concentrating on FP and FN rates, we enable a more granular analysis of error distributions and consumer preferences.
Unlike Feng et al. (2022), where firms optimize directly over bias and variance-based errors, the FP and FN rates in our model simultaneously determine both the first and second moments of the cost function. Consumers care about the expected cost of a mistake, and a lower expected cost does not necessarily imply a higher variance. This approach is particularly effective for modeling consumer behavior, making it easier to see how different types of mistakes affect consumer satisfaction.
@@
Third, we introduce heterogeneity among both consumers and firms. Empirical evidence suggests that consumers exhibit varying sensitivities to FP and FN errors (Cappelen et al., 2023). We formally model this heterogeneity using FP and FN cost parameters, accounting for cardinal differences, the magnitude of costs associated with FP and FN errors for each consumer, and ordinal differences, the relative importance or priority that consumers assign to FP versus FN errors. Similarly, we incorporate firm heterogeneity through varying error endowments. Based on their initial endowments, firms may specialize in optimizing for different error types or target specific consumer segments, leading to potentially diverse equilibria.
@@ 3.1 Perfectly Positively Correlated Cost Sensitivities
We now impose structure on consumers’ cost sensitivities, first considering the case where $\alpha=\beta$, implying perfectly positively correlated consumer sensitivities to FP and FN rates. Under this structure, the model simplifies to the Hotelling model (Hotelling, 1981; d’Aspremont et al., 1979). For conciseness, we define $\mathcal{F}_{j}$ as the total error rates for firm $j$: $\mathcal{F}_{j}=FP_{j}+FN_{j}$. The consumer’s cost function $c_{i}(j)$ then simplifies to:
@@ 4.3 Investment Costs
Much of the artificial intelligence and machine learning literature on scaling laws (Kaplan et al., 2020; Hoffmann et al., 2022) has demonstrated that dataset size and performance/error are highly intertwined. In our framework, this implies that the effectiveness of firm investments are directly related to their latent dataset size variable. We propose that the training dataset size each firm has directly effects its investment costs: a firm with a much larger training dataset will be able to invest more efficiently on improving their error rates as opposed to a firm with much fewer training samples.
Marginal costs should also be different across FP and FN rates. There should a distinct tradeoff for each firm on reducing their FP and FN rates (Tirole, 1988; Mas-Colell et al., 1995). In particular, if a firm already has a low FP error rate, it should be more expensive to lower it even further. A firm with $FP=0.1$ and investing to reduce it to 0.09 should experience much higher costs than a firm attempting to lower its FP rate from 0.3 to 0.29.

### v3

@@
Keywords: AI model market, model error trade-off, heterogeneous cost sensitivities, horizontal differentiation, Hotelling’s model.
@@ 1 Introduction
While this goal is theoretically optimal, achieving a minimal error rate can be both costly and impractical for firms (pmlr-v97-zhang19p). Improving multiple error dimensions simultaneously can be expensive and inefficient. In addition, an improvement in one dimension may lead to a worsening in another dimension, inducing an inverse trade-off across dimensions (yaniv1995graininess). A typical example of this is binary classification, where directly minimizing both Type I and Type II errors with a fixed sample size is typically infeasible (mudge2012setting). Assuming a fixed total level of error, designing a model with fewer false positive (FP) errors would shrink the rejection region but also simultaneously lead to an increase in false negative (FN) error risk, and vice versa. As a result, model designers have to prioritize between FP and FN errors based on their expectations about consumer preferences. A standard statistical approach is fixing one dimension and optimizing the other (doi:10.1177/25152459221080396).
On the demand side, consumers often care less about the total model error and more about the specific types of errors a model makes, as some errors may be more costly to them than others. The bias-variance trade-off is an example of this inverse relation in model performance; lowering bias by improving model complexity typically leads to higher variance (geman1992neural). One specific measure of model performance that consumers often consider is based on FP and FN rates (cappelen2023second). Different consumers exhibit varying levels of sensitivities to FP and FN rates, depending on their specific use cases and their associated cost sensitivities to FP and FN errors.
A prevalent shortcoming with generative AI models is hallucination, where the model generates entirely incorrect outputs. Hallucination rates, or more broadly FP rates, of the most popular models like ChatGPT remain quite significant (rawte2023surveyhallucinationlargefoundation). Users employing AI for creative writing may have a completely different tolerance for hallucination compared to someone using the same model for disease diagnosis. This consumer heterogeneity can be captured by different sensitivities toward hallucination rates. For example, a scientific user may prefer a non-informative model output (i.e., “I don’t know”) as opposed to a model hallucination.
@@ 2 Related Literature
Our work builds upon past research at the intersection of machine learning and strategic behavior, focusing on how competition affects efficiency and the resulting equilibria. Previous literature has extensively analyzed competition in such settings. For example, 10.1145/1993636.1993666 propose an optimization-based framework for strategic competition, but does not include prices. feng2022bias examine the bias-variance trade-off in competitive environments, demonstrating a tendency for firms to favor variance-based errors over bias-based errors. Similarly, jagadeesan2024improvedbayesriskyield propose a classification model under which competition can potentially lead to worse overall model accuracy. Other studies adopt more information-driven approaches to competition (liang2020gamesincompleteinformationplayed) and explore coordination among firms to learn about consumer types (gradwohl2023coopetition).
We contribute to the long-standing industrial organization literature by presenting a model of horizontal differentiation in an AI model market. Consumers’ cost sensitivities govern their purchasing behavior of models and thus determine market equilibria. shaked1982relaxing present a multi-stage game model with vertical differentiation, but assume consumers have identical tastes. gabszewicz1979price also assume no horizontal differentiation but discuss the notion of firm specialization.
This paper differs from the related literature along multiple dimensions. First, we focus explicitly on competition over false positive (FP) and false negative (FN) rates rather than on the bias-variance trade-off. While the bias-variance framework is instrumental for understanding model complexity and generalization (feng2022bias), the FP/FN trade-off is more specific to classification tasks. jagadeesan2024improvedbayesriskyield examines competition based on accuracy measured with Bayes risk, but firms do not choose prices. Classification problems are central to many real-world applications, such as medical diagnostics, fraud detection, and content filtering, where FP and FN errors carry distinct and significant costs. By concentrating on FP and FN rates, we enable a more granular analysis of error distributions and consumer preferences.
Unlike feng2022bias, where firms optimize directly over bias and variance-based errors, the FP and FN rates in our model simultaneously determine both the first and second moments of the cost function. Consumers care about the expected cost of a mistake, and a lower expected cost does not necessarily imply a higher variance. This approach is particularly effective for modeling consumer behavior, making it easier to see how different types of mistakes affect consumer satisfaction.
@@
Third, we introduce heterogeneity among both consumers and firms. Empirical evidence suggests that consumers exhibit varying sensitivities to FP and FN errors (cappelen2023second). We formally model this heterogeneity using FP and FN cost parameters, accounting for cardinal differences, the magnitude of costs associated with FP and FN errors for each consumer, and ordinal differences, the relative importance or priority that consumers assign to FP versus FN errors. Similarly, we incorporate firm heterogeneity through varying error endowments. Based on their initial endowments, firms may specialize in optimizing for different error types or target specific consumer segments, leading to potentially diverse equilibria.
@@ 3.1 Perfectly Positively Correlated Cost Sensitivities
We now impose structure on consumers’ cost sensitivities, first considering the case where $\alpha=\beta$, implying perfectly positively correlated consumer sensitivities to FP and FN rates. Under this structure, the model simplifies to the Hotelling model (hotelling1981stability; d1979hotelling). For conciseness, we define $\mathcal{F}_{j}$ as the total error rates for firm $j$: $\mathcal{F}_{j}=FP_{j}+FN_{j}$. The consumer’s cost function $c_{i}(j)$ then simplifies to:
@@ 4.3 Investment Costs
Much of the artificial intelligence and machine learning literature on scaling laws (kaplan2020scaling; hoffmann2022training) has demonstrated that dataset size and performance/error are highly intertwined. In our framework, this implies that the effectiveness of firm investments are directly related to their latent dataset size variable. We propose that the training dataset size each firm has directly effects its investment costs: a firm with a much larger training dataset will be able to invest more efficiently on improving their error rates as opposed to a firm with much fewer training samples.
Marginal costs should also be different across FP and FN rates. There should a distinct tradeoff for each firm on reducing their FP and FN rates (tirole1988theory; mas1995microeconomic). In particular, if a firm already has a low FP error rate, it should be more expensive to lower it even further. A firm with $FP=0.1$ and investing to reduce it to 0.09 should experience much higher costs than a firm attempting to lower its FP rate from 0.3 to 0.29.